OpenStudy (anonymous):

arithmetic series find the correct sum 7+14+21...+98

OpenStudy (anonymous):

answers 735 -25/3 -245/6

OpenStudy (anonymous):

I guess you can do it girl.. In this series, first term a=7, last term l= 98 , common difference d = 7 You see that the sum of series will be, n/2 (a+l)... We have to find n now, since 'l' is the last or nth term l = a + (n-1). d or, 98 = 7 + (n-1) * 7 or, 91 = 7 (n-1) or, n-1 = 13 so, n= 14 Hence, Sum= n/2 (a+l) = 14/2 (7+98) = 7 * 105 = 735

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem.. anytime here to help